OpenStudy (chosenmatt):

Which line shows the first error in the solution?(1)3(x-2)=18 (2)3x-6=18 (3)3x-6+6=18+6 (4)3x=24 (5)x=21

OpenStudy (chosenmatt):

@TheSmartOne

OpenStudy (anonymous):

need help

OpenStudy (chosenmatt):

me too XD

OpenStudy (anonymous):

I ask pepele to help I dot no

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

my bruther I tring to help

OpenStudy (chosenmatt):

k thx :(

OpenStudy (anonymous):

give him a sec

OpenStudy (anonymous):

try this website :Linear Equation Calculator

OpenStudy (anonymous):

try it

OpenStudy (anonymous):

:-)

OpenStudy (chosenmatt):

I can't get to it :(

OpenStudy (anonymous):

did you typeing it on goole

OpenStudy (anonymous):

tipe it in to google

OpenStudy (anonymous):

hi whats the question

OpenStudy (chosenmatt):

I STILL DON'T GET IT THIS IS A NIGHTMARE LOL I BEEN ON THIS QUESTION FOOR ABOUT 20MINS!!! LOL XD

OpenStudy (anonymous):

wow lol whats the question

OpenStudy (anonymous):

x=9

OpenStudy (chosenmatt):

Which line shows the first error in the solution?(1)3(x-2)=18 (2)3x-6=18 (3)3x-6+6=18+6 (4)3x=24 (5)x=21

OpenStudy (anonymous):

there is no error

OpenStudy (chosenmatt):

yes there is I just don't know which one

OpenStudy (anonymous):

wow hard question

OpenStudy (chosenmatt):

@StudyGurl14

OpenStudy (anonymous):

@chosenmatt can you help me in a queston plzz

OpenStudy (studygurl14):

Can you retype the question so that each part is on a seperate line? It is too hard to read when smashed together like that @chosenmatt

OpenStudy (chosenmatt):

@StudyGurl14 number with the parantheses is the lines

OpenStudy (anonymous):

@StudyGurl14 or @chosenmatt can you help me in this question plzz Identify the type of sample used. Max surveys the children who live on his street about the types of equipment they want installed at the new city playground. A. convenience B. simple random C. stratified random D. systematic

OpenStudy (studygurl14):

@lovepurple DO NOT post your question in someone else's question. It is rude and against the code of conduct.

OpenStudy (studygurl14):

(1)3(x-2)=18 (2)3x-6=18 (3)3x-6+6=18+6 (4)3x=24 (5)x=21 Right @chosenmatt ?

OpenStudy (anonymous):

it is sorry

OpenStudy (studygurl14):

It's ok @lovepurple . Just don't do it again.

OpenStudy (chosenmatt):

YUP!! and it's okay @lovepurple

OpenStudy (anonymous):

so can you help me???

OpenStudy (studygurl14):

@lovepurple post it in your own thread and link me. I'll get to it when i'm done here

OpenStudy (anonymous):

ok tnx @StudyGurl14

OpenStudy (chosenmatt):

ill help too @lovepurple

OpenStudy (studygurl14):

The mistake is the last line. Instead of dividing 24 by 3, the "person" subtracted 3 from 24 to get 21

OpenStudy (chosenmatt):

okay can you help me with 4 more?

OpenStudy (studygurl14):

not in this thread. one question per thread.

OpenStudy (chosenmatt):

okay

OpenStudy (studygurl14):

two is ok, but this thread is really long already. too long of threads make for lagging and loading problems

OpenStudy (studygurl14):

can i have a medal please? @chosenmatt

OpenStudy (chosenmatt):

so it's x=21 @StudyGurl14

OpenStudy (studygurl14):

yea. line 5